# Optimal Liouville theorems for the Lane-Emden equation on Riemannian manifolds

### v1

@@ Abstract
We study degenerate quasilinear elliptic equations on Riemannian manifolds and obtain several Liouville theorems. Notably, we provide a rigorous proof asserting the nonexistence of positive solutions to the subcritical Lane-Emden-Fowler equations over complete Riemannian manifolds with nonnegative Ricci curvature. These findings serve as a significant generalization of Gidas and Spruck’s pivotal work (Comm. Pure Appl. Math. 34, 525-598, 1981) which focused on the semilinear case, as well as Serrin and Zou’s contributions (Acta Math. 189, 79-142, 2002) within the context of Euclidean geometries.
@@ 1 Introduction
In the seminal paper [15] Gidas and Spruck proved that there are no positive solutions to the Lane-Emden equation (see [15, Theorem 1.1]), i.e.,
@@
$1<\alpha<\dfrac{n+2}{n-2}.$
This result can be regarded as an extension of the classical Liouville theorem for harmonic functions defined on Euclidean space states that any real, positive, and entire harmonic function must be a constant.
Gidas and Spruck’s result can be generalized to quasilinear elliptic equations. For example, Serrin and Zou [41, Corollary II] proved that the classical Lane-Emden-Fowler equation
@@
The equations (1.1) and (1.2) on an Euclidean space $\mathbb{R}^{n}$ have been extensively studied by many mathematicians. Without being exhaustive with the huge amount of references concerned with this topic, let us mention the works [3, 5, 6, 8, 37, 31, 34] and references therein. Besides, in [15, 41] the following equation
@@
was also be studied and some universal estimates for positive solutions to (1.3) were established (also see [13, 39, 40]). In particular, Yanyan Li and Lei Zhang [30] in 2003 studied the equation and proved that the following results [30, Theorem 1.3]: Suppose that $f$ is locally bounded on $(0,+\infty)$ and $s^{-\frac{n+2}{n-2}}f(s)$ is non-increasing on $(0,+\infty)$. Assume that $u$ is a positive classical solution of the above equation (1.3) in $\mathbb{R}^{n}$ with $n\geq 3$, then either for some $b>0$,
@@
$-\Delta u+\dfrac{n-2}{4(n-1)}Ru=\lambda u^{\frac{n+2}{n-2}}$
@@
In fact, whether or not Liouville theorems pertaining to the Lane-Emden-Fowler equation on Riemannian manifolds with nonnegative Ricci curvature are true for
$-\infty<\alpha<\frac{(n+1)p-n}{n-p}\quad\mbox{and}\quad 1<p<n$
is a long-standing open problem.
It should be mentioned that Gidas and Spruck [15, Theorem 1.2] demonstrated that no positive solution to the Lane-Emden equation (1.1) exists on any complete Riemannian manifold of dimension $n$ with nonnegative Ricci curvature, under the condition
$1<\alpha<\dfrac{n+2}{n-2}.$
@@
The existence of such kind of cutoff functions on complete Riemannian manifolds is highly nontrivial. If the Ricci curvature is nonneagtive, then one can prove the existence of Laplacian cutoff functions from Riemannian rigidity theory (cf. [7, Theorem 6.33] or [47, Lemma 1.5]), with a careful scaling argument (see [18, Theorem 2.2, (b)] or [2, Corollary 2.3]). On the other hand, very recently the first and third named author and Wei [23] have shown that (1.2) does not admits any positive solution if $p>1$ and
@@
The goal of this paper is to answer the above problem, i.e., to show the optimal Liouville theorem and establish some universal $\log$-gradient estimates for Lane-Emden-Fowler equation equation. More concretely, we will show that the equation $\Delta_{p}u+u^{\alpha}=0$ does not admit any positive solution if $\alpha\in(-\infty,\,p_{s})$ where $p_{s}$ is the Sobolev critical exponent given by
$p_{s}\equiv\dfrac{(n+1)p-n}{(n-p)^{+}}$
(see Theorem 2.1); and we also establish an universal $\log$-gradient estimate for positive solutions to (1.3) on a complete manifold (see Theorem 2.4 and Theorem 2.5).
Here, by using the word “universal” here, we mean that our bounds are not only independent of any given solution under consideration but also do not require, or assume, any boundary conditions whatsoever. For instance, the exact Cheng-Yau $\log$-gradient estimates is an universal a priori gradient estimates.
@@
Using Cheng-Yau’s method, Li [29] considered the more general Lane-Emden type equation (1.1) and obtained several gradient estimates and Harnack inequalities on Riemannian manifolds with nonnegative Ricci curvature under the assumption $1\leq\alpha<\frac{n}{n-2},\ n\geq 4$. Notably, he established a Liouville theorem concerning the nonexistence of positive solutions. This result was further extended by Ma, Huang, and Luo [33] when $n\geq 5$. Inspired by Wang and Zhang’s approach, the third-named author and Wei [52] employed the Moser iteration strategy to prove a local gradient estimate of the Cheng-Yau type for positive solutions to the semilinear elliptic equation (1.1) on Riemannian manifolds. Specifically, they derived a Liouville theorem for the Lane-Emden equation on complete Riemannian manifolds with dimension $n$ and nonnegative Ricci curvature, given that
@@
This finding was later refined by the first and third named author and Wei [23], who expanded the range of $\alpha$ to
@@
where $h$ is a smooth function, and employ the Nash-Moser iteration technique to obtain some refined gradient estimates of the solutions to the above equation, if $(M^{n},\,g)$ is of Ricci curvature lower bound and $h$ fulfills some suitable assumptions. This methodology can be applied to address other problems as well. Additional related results can be found in various sources [49, 26, 19, 54, 48, 55, 54, 51, 22, 20, 21, 24].
@@
Grigor’yan and Sun [16] examined the uniqueness of a nonnegative solution to this inequality for $p=2$. Notably, they utilized the condition of volume growth instead of relying on nonnegative Ricci curvature. Similar results hold for general $p>1$ [42]. That is, there exists a non-trivial positive solution to the inequality (1.4) provided $\alpha>p-1$ and for some fixed point $o\in M^{n}$ and some positive number $C$
@@
This provides a new proof of Bidaut-Veron and Pohozaev’s result [4, Theorems 3.3 (iii) and 3.4 (ii)] for the Euclidean case when $1<p<n$. It should be mentioned that the range of $\alpha$ can be extended to
@@
by Serrin and Zou [41, Theorem II (d)]. Other related results can be found in various sources [32, 43, 55, 17, 53, 57, 25] and references therein.
@@ 2 Main results and ideas of proof
In the original definition of subcritical by Serrin and Zou [41, (1.7)], $f$ is required to be non-negative and the constant $\alpha$ in (2.2) is required to lie in $\left(0,\frac{np}{n-p}-1\right)$. In our theorem, we do not require this constraints. In section 4, $f$ can be negative and $\alpha$ can also be negative.
@@
Let $M^{n}$ be a complete and $n$-dimensional Riemannian manifold with the Ricci curvature bounded from below by a nonpositive constant $-(n-1)\kappa$. Assume for some constant $\delta_{0}<1$,
@@
If $p>1$ and $u$ is a positive and weak solution to (2.1), then there exists a positive constant $C_{n,p,\delta_{0}^{+}}$ depending only on $n,p$ and $\delta_{0}^{+}$ such that
@@
Let $M^{n}$ be a complete and $n$-dimensional Riemannian manifold with the Ricci curvature bounded from below by a nonpositive constant $-(n-1)\kappa$. Assume $p>1$ and the condition (2.3) holds. If $u$ is a positive and weak solution to (2.1), then
@@
in the Euclidean space. However, it is recognized that the lower bound of the Ricci curvature generally does not constrain the Hessian of the distance function, in other words, such a cutoff function $\zeta$ appears non existent for any $R>0$ in general Riemannian manifolds characterized by nonnegative Ricci curvature. So we can not use the method of Gidas and Spruck [15] or Serrin and Zou [41] directly. Fortunately, to obtain the Liouville theorem, one does not need any information of the Hessian of the distance function. However, one needs the information of the lower bound of the Ricci curvature to control the growth of volume.
@@
As an immediately consequence, we have the following generalization of Serrin and Zou [41, Theorem II (b, c)] when $M^{n}=\mathbb{R}^{n}$ and $1<p<n$.
@@
where $q>1$ and $\lambda>0$, and provide a new and simple proof of Bidaut-Veron and Veron’s Liouville result [5, Theorem 6.1] on a compact manifold (see Theorem 5.2).
@@ 3 Basic formulas and inequalities
According to the classical regularity theory (for example, see [14, 45, 46]), we know that
$u\in C_{loc}^{1,\beta}(\Omega)\cap W^{2,2}_{loc}(\Omega)\quad\mbox{and}\quad u\in C_{loc}^{2}(\Omega^{c}_{cr}).$
@@
It is worthy to point out that the endomorphisms $\mathcal{U}$ and $\mathcal{X}$ may have singularities on $\Omega_{cr}$. But we have the following conclusions on regularity.
@@
$\mathcal{U},\quad\mathcal{X}\in L^{2}_{loc}(\Omega).$
@@
When $p\geq 2$, we have $\mathcal{U},\mathcal{X}\in L^{2}_{loc}(\Omega)$ (for example see [41, Corollary in P. 117]). Now we consider the case $21<p<$ 2. According to [41, Proposition 8.1], for any compact set $K$,
@@
$\mathcal{U},\quad\mathcal{X}\in L^{2}_{loc}(\Omega).$
@@
$\operatorname{div}(\mathbf{X})=bu^{b-1}\left\lvert\nabla u\right\rvert^{p}-u^{b}f(u)\in C_{loc}^{0}(\Omega),$
thus $\mathscr{X}\in L^{2}_{loc}(\Omega)$. By the Lemma due to Serrin and Zou [41, Lemma 6.3], we know that
@@
$\mathrm{tr}\left(\mathscr{X}^{2}\right)=\left(\mathcal{X}_{i}^{j}-\frac{1}{n}\mathrm{tr}\mathcal{X}\delta_{i}^{j}\right)\left(\mathcal{X}^{i}_{j}-\frac{1}{n}\mathrm{tr}\mathcal{X}\delta_{j}^{i}\right)=\mathrm{tr}(\mathcal{X}^{2})-\frac{1}{n}(\mathrm{tr}\mathcal{X})^{2}$
@@
$\left\lvert\mathscr{X}\right\rvert^{2}=\left(\mathcal{X}_{i}^{j}-\frac{\mathrm{tr}\mathcal{X}}{n}\delta_{i}^{j}\right)\left(\mathcal{X}_{i}^{j}-\frac{\mathrm{tr}\mathcal{X}}{n}\delta_{i}^{j}\right)=\mathrm{tr}\left(\mathcal{X}\mathcal{X}^{T}\right)-\frac{1}{n}(\mathrm{tr}\mathcal{X})^{2}.$
@@
On the one hand, using the fact that for every smooth function $\phi$ and tangent vector fields $Y$ and $Z$,
$\nabla_{\phi Y}Z-\operatorname{div}\left(\phi Y\right)Z=\nabla_{Y}\left(\phi Z\right)-\operatorname{div}Y\left(\phi Z\right)=\phi\left(\nabla_{Y}Z-\operatorname{div}YZ\right),$ (3.6)
we compute from (3.2) and (3.6)
@@
We have the following fundamental and important pointwise identity.
@@
$\begin{split}&\operatorname{div}\left(u^{a}\left(\mathscr{X}(\mathbf{X})\right)-\dfrac{n(p-1)a+(n-1)pb}{np}u^{a-1}\left\langle\mathbf{X},\nabla u\right\rangle\mathbf{X}\right)\\ =&u^{a}\left(\mathrm{tr}(\mathscr{X}^{2})+\mathrm{Ric}\left(\mathbf{X},\mathbf{X}\right)\right)\\ &+\left(\dfrac{(p-1)a}{p}\left(1-\dfrac{(n-p)a}{(n-1)p}\right)-\dfrac{n-1}{n}\left(b+\dfrac{n(p-1)a}{(n-1)p}\right)^{2}\right)u^{a+2b-2}\left\lvert\nabla u\right\rvert^{2p}\\ &+\dfrac{n-1}{n}\left(\dfrac{((n+1)p-n)a}{(n-1)p}f(u)-uf^{\prime}(u)\right)u^{a+2b-1}\left\lvert\nabla u\right\rvert^{p}.\end{split}$ (3.9)
@@
$\begin{split}&\operatorname{div}\left(u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\left(\mathscr{X}(\mathbf{X})\right)\right)\\ =\,&u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\left(\mathrm{tr}(\mathscr{X}^{2})+\mathrm{Ric}\left(\mathbf{X},\mathbf{X}\right)\right)\\ &+\dfrac{(n-1)(p-1)\alpha}{(n+1)p-n}\left(1-\dfrac{(n-p)\alpha}{(n+1)p-n}\right)u^{-\frac{((n+1)p-2n)\alpha}{(n+1)p-n}-2}\left\lvert\nabla u\right\rvert^{2p}\\ &+\dfrac{n-1}{n}\left(\alpha f(u)-uf^{\prime}(u)\right)u^{-\frac{((n+1)p-2n)\alpha}{(n+1)p-n}-1}\left\lvert\nabla u\right\rvert^{p}.\end{split}$ (3.10)
@@
$=$ $u^{a}\left(\mathrm{tr}(\mathscr{X}^{2})+\mathrm{Ric}\left(\mathbf{X},\mathbf{X}\right)\right)$
@@
=\,= $\operatorname{div}\left(u^{a+2b}\Delta_{p}u\mathbf{U}\right)-\left\langle\nabla\left(u^{a+2b}\Delta_{p}u\right),\mathbf{U}\right\rangle$
@@
=\,= $u^{a}\left(\mathrm{tr}(\mathscr{X}^{2})+\mathrm{Ric}\left(\mathbf{X},\mathbf{X}\right)\right)$
@@
we get the desired formula (3.10) and complete the proof formally. By 3.2, we know that $\mathrm{tr}\left(\mathcal{X}^{2}\right)\in L^{1}_{loc}$. It follows that every term on the right hand side of 3.5 lies in $L^{1}_{loc}$. ∎
In the end of this section, we collect some useful inequalities in geometric analysis. Let $M^{n}$ be a Riemannian manifold with Ricci curvature bounded from below by some nonnegative constant $-(n-1)\kappa$. The first one is the following Saloff-Coste’s Dirichlet Sobolev inequality: for every $\psi\in C_{0}^{\infty}\left(B_{R}\right)$
$\left(\fint_{B_{R}}\left\lvert\psi\right\rvert^{\frac{n}{n-1}}\right)^{\frac{n-1}{n}}\leq e^{C_{n}\left(1+\sqrt{\kappa}R\right)}R\left(\fint_{B_{R}}\left\lvert\nabla\psi\right\rvert+R^{-1}\fint_{B_{R}}\left\lvert\psi\right\rvert\right).$

### v3

@@ Abstract
This paper study the Liouville-type theorems and universal estimates of positive solutions for the $p$-Laplacian Lane-Emden type equations on Riemannian manifolds. In particular, we prove an optimal Liouville theorem, demonstrating the nonexistence of positive solutions for subcritical $p$-Laplacian Lane-Emden equations on complete noncompact Riemannian manifolds with nonnegative Ricci curvature. This resolves a longstanding open problem in the field. Our approach features two key innovations: (i) an adapted Nash-Moser iteration scheme yielding universal log-gradient estimates for the degenerate elliptic operator, and (ii) new integral estimates that produce sharp nonexistence results. These techniques enable us to extend the Euclidean quasilinear theory of Serrin-Zou (Acta Math. 189, 79-142, 2002) to general Riemannian manifolds.
@@ 1 Introduction
In the seminal paper [22] Gidas and Spruck proved that there are no positive solutions to the Lane-Emden equation (see [22, Theorem 1.1]), i.e.,
@@
$1\leq\alpha<\dfrac{n+2}{n-2}.$
This result can be regarded as an extension of the classical Liouville theorem for harmonic functions defined on the Euclidean space states that any real, positive, and entire harmonic function must be a constant.
Gidas and Spruck’s result can be generalized to quasilinear elliptic equations. For example, Serrin and Zou [52, Corollary II] proved that the classical Lane-Emden-Fowler equation
@@
The equations (1.1) and (1.2) on an Euclidean space $\mathbb{R}^{n}$ have been extensively studied by many mathematicians. Without being exhaustive with the huge amount of references concerned with this topic, let us mention the works [4, 6, 7, 11, 21, 46, 39, 43, 49] and references therein. Besides, in [22, 52] the following equation
@@
was also be studied and some universal estimates for positive solutions to (1.3) were established (also see [17, 50, 51]). In particular, Yanyan Li and Lei Zhang [38] in 2003 studied the equation and proved the following results [38, Theorem 1.3]: Suppose that $f$ is locally bounded on $(0,+\infty)$ and $s^{-\frac{n+2}{n-2}}f(s)$ is non-increasing on $(0,+\infty)$. Assume that $u$ is a positive classical solution of the above equation (1.3) in $\mathbb{R}^{n}$ with $n\geq 3$, then either for some $b>0$,
@@
$-\Delta_{g}u+\dfrac{n-2}{4(n-1)}Ru=\lambda u^{\frac{n+2}{n-2}}$
@@
In fact, whether or not Liouville theorems pertaining to the Lane-Emden-Fowler equation on Riemannian manifolds with nonnegative Ricci curvature are true for -¡α¡(n+1)p-$n$-p and  1¡p¡n is a long-standing open problem.
It should be mentioned that Gidas and Spruck [22, Theorem 1.2] demonstrated that no positive solution to the Lane-Emden equation (1.1) exists on any complete Riemannian manifold of dimension $n$ with nonnegative Ricci curvature, under the condition
$1\leq\alpha<\dfrac{n+2}{n-2}.$
@@
It is highly nontrivial for us to prove the existence of such kind of cutoff functions on complete Riemannian manifolds. Indeed, if the Ricci curvature is nonneagtive, then one can prove the existence of Laplacian cutoff functions from Riemannian rigidity theory (cf. [10, Theorem 6.33] or [60, Lemma 1.5]), with a careful scaling argument (see [25, Theorem 2.2, (b)] or [3, Corollary 2.3]). Recently, Catino and Monticelli [8] gave a new proof of the Liouville theorem [22, Theorem 1.2] on the equation (1.1). In fact, Catino and Monticelli [8, Theorem 1.7] made also use of the same cutoff function as in [22, Theorem 1.2] to prove the conclusion. Later on Lu [41] used maximum priciple to establish Cheng-Yau type gradient estimate and hence to prove the Liouville theorem due to Gidas and Spruck. Very recently, Ciraolo, Farina, and Polvara employed a different method to provide a very simple proof in [13].
On the other hand, very recently the first and third named author and Wei [31] have shown that (1.2) does not admits any positive solution if $p>1$ and
@@
The goal of this paper is to answer the above problem, i.e., to show the optimal Liouville theorem and establish some universal $\log$-gradient estimates for Lane-Emden-Fowler equation. More concretely, we will show that the equation $\Delta_{p}u+u^{\alpha}=0$ does not admit any positive solution if $\alpha\in(-\infty,\,p_{s})$ where $p_{s}$ is the Sobolev critical exponent given by p_s(n+1)p-n(n-p)+ (see Theorem 2.1); and we also establish an universal $\log$-gradient estimate for positive solutions to (1.3) on a complete manifold (see Theorem 2.4 and Theorem 2.5).
Here, by using the word “universal”, we mean that our bounds are not only independent of any given solution under consideration but also do not require, or assume, any boundary conditions whatsoever. For instance, the exact Cheng-Yau $\log$-gradient estimates is an universal a priori gradient estimates.
@@
Using Cheng-Yau’s method, Li [37] considered the more general Lane-Emden type equation (1.1) and obtained several gradient estimates and Harnack inequalities on Riemannian manifolds with nonnegative Ricci curvature under the assumption $1\leq\alpha<\frac{n}{n-2},\ n\geq 4$. Notably, he established a Liouville theorem concerning the nonexistence of positive solutions. When $n\geq 5$, Ma, Huang, and Luo [42] proved a gradient estimate by which one can derive the nonexistence of positive solutions for
$\infty<\alpha<\frac{2n^{2}+9n+6}{2n(n+2)}.$
The De Giorgi-Nash-Moser iteration method holds significant influence not only in the analysis of PDEs but also in geometric analysis. Zhang and Zhu [70] introduced a variant of the Bochner formula applicable to Alexandrov spaces. By combining this adapted formula with the Nash-Moser iteration method, they derived Yau’s gradient estimate for harmonic functions in these settings. More recently, Bellettini [2] utilized the De Giorgi iteration to provide an alternative proof for the Schoen-Simon-Yau curvature estimates and related Bernstein-type theorems, further extending the original findings to encompass the case of six-dimensional stable minimal immersions.
Inspired by Wang and Zhang’s approach, the third-named author of this paper and Wei [64] employed the Moser iteration strategy to prove a local gradient estimate of the Cheng-Yau type for positive solutions to the semilinear elliptic equation (1.1) on Riemannian manifolds. Specifically, they derived a Liouville theorem for the Lane-Emden equation on complete Riemannian manifolds with dimension $n$ and nonnegative Ricci curvature, given that
@@
This finding was later refined by the first and third named author of this paper and Wei [31], who expanded the range of $\alpha$ to
@@
where $h$ is a smooth function, and employ the Nash-Moser iteration technique to obtain some refined gradient estimates of the solutions to the above equation, if $(M^{n},\,g)$ is of Ricci curvature lower bound and $h$ fulfills some suitable assumptions. This methodology can be applied to address other problems as well. Additional related results can be found in various sources [26, 27, 28, 29, 30, 31, 32, 34, 62, 61, 66].
@@
Grigor’yan and Sun [23] examined the uniqueness of a nonnegative solution to this inequality for $p=2$. Notably, they utilized the condition of volume growth instead of relying on nonnegative Ricci curvature. Similar results hold for general $p>1$ [54]. That is, there exists a no positive solution to the inequality (1.4) provided $\alpha>p-1$ and for some fixed point $o\in M^{n}$ and some positive number $C$
@@
This provides a new proof of Bidaut-Veron and Pohozaev’s result [5, Theorems 3.3 (iii) and 3.4 (ii)] for the Euclidean case when $1<p<n$. It should be mentioned that the range of $\alpha$ can be extended to
@@
by Serrin and Zou [52, Theorem II (d)]. Other related results can be found in various sources [33, 40, 55, 66, 24, 65, 69] and references therein.
Besides, some mathematicians studied the classification to positive solutions to critical $p$-Laplace equation on a complete Riemannian manifold, which is just equation (1.1) with $\alpha=p_{s}$, i.e.,
$\Delta_{p}u+u^{p_{s}}=0.$
For more details we refer to [8, 9, 13, 48, 59] and references therein. We also should mentioned that one studied the Liouville equation which is closely related to Lane-Emden equation and Yamabe equation, we refer to [19, 20] and references therein.
@@ 2 Main results and ideas of proof
Our attention is directed toward the subcritical case. A subsequent paper [53] will address the critical case.
@@
In the original definition of subcritical by Serrin and Zou [52, (1.7)], $f$ is required to be non-negative and the constant $\alpha$ in (2.2) is required to lie in $\left(0,\,\frac{np}{n-p}-1\right)$. In our theorem, we do not require this constraints. In section 4, $f$ can be negative and $\alpha$ can also be negative.
@@
Let $M^{n}$ be a complete and $n$-dimensional Riemannian manifold with the Ricci curvature bounded from below by a nonpositive constant $(n-1)\kappa$. Assume for some constant $\delta_{0}<1$,
@@
If $p>1$ and $u$ is a positive and weak solution to (2.1), then there exists a positive constant $C_{n,p,\delta_{0}^{+}}$ depending only on $n,p$ and $\delta_{0}^{+}=\max\{0,\delta_{0}\}\in[0,\,1)$ such that
@@
Let $M^{n}$ be a complete and $n$-dimensional Riemannian manifold with the Ricci curvature bounded from below by a nonpositive constant $(n-1)\kappa$. Assume $p>1$ and the condition (2.3) holds. If $u$ is a positive and weak solution to (2.1), then
@@
in the Euclidean space. However, it is recognized that the lower bound of the Ricci curvature generally does not constrain the Hessian of the distance function, in other words, such a cutoff function $\zeta$ appears non existent for any $R>0$ in general Riemannian manifolds characterized by nonnegative Ricci curvature. So we can not use the method of Gidas and Spruck [22] or Serrin and Zou [52] directly. Fortunately, to obtain the Liouville theorem, one does not need any information of the Hessian of the distance function. However, one needs the information of the lower bound of the Ricci curvature to control the growth of volume. In the semilinear case where $p=2$, Ciraolo, Farina, and Polva [13, Theorem 1.4] effectively employed the distance cutoff function to provide an alternative proof of the findings by Gidas and Spruck [22, Theorem 1.2].
@@
As an immediately consequence, we have the following generalization of Serrin and Zou [52, Theorem II (b, c)] when $M^{n}=\mathbb{R}^{n}$ and $1<p<n$.
@@
where $q>1$ and $\lambda>0$, and provide a new and simple proof of Bidaut-Veron and Veron’s Liouville result [6, Theorem 6.1] on a compact manifold (see Theorem 5.2).
@@ 3 Basic formulas and inequalities
According to Theorem 1.4 in [1] and the classical regularity theory (for example, see [18, 57, 58]), we know that
$u\in C_{loc}^{1,\beta}(\Omega)\cap W^{2,2}_{loc}(\Omega\setminus\Omega_{cr})\quad\mbox{and}\quad u\in C_{loc}^{\infty}(\Omega^{c}_{cr}).$
@@
It is worthy to point out that the endomorphisms $\mathcal{U}$ and $\mathcal{X}$ may have singularities on $\Omega_{cr}$. But we have the following conclusions on regularity which is implied in [1]. For the sake of convenience, we give a proof here.
@@
$\mathcal{U}~~\mbox{and}~~\mathcal{X}\in L^{2}_{loc}(\Omega).$
@@
When $p\geq 2$, we have $\mathcal{U},\mathcal{X}\in L^{2}_{loc}(\Omega)$ (for example see [52, Corollary in P. 117]). Now we consider the case $1<p<2$. According to [52, Proposition 8.1], for any compact set $K$,
@@
$\mathcal{U}~~\mbox{and}~~\mathcal{X}\in L^{2}_{loc}(\Omega).$
Thus, we finish the proof. ∎
@@
$\operatorname{div}\mathbf{X}=bu^{b-1}\left\lvert\nabla u\right\rvert^{p}-u^{b}f(u)\in C_{loc}^{0}(\Omega),$
thus $\mathscr{X}\in L^{2}_{loc}(\Omega)$. By the Lemma due to Serrin and Zou [52, Lemma 6.3], we know that
@@
$\mathrm{tr}\left(\mathscr{X}^{2}\right)=\left(\mathcal{X}_{i}^{j}-\frac{1}{n}\mathrm{tr}\mathcal{X}\delta_{i}^{j}\right)\left(\mathcal{X}^{i}_{j}-\frac{1}{n}\mathrm{tr}\mathcal{X}\delta_{j}^{i}\right)=\mathrm{tr}(\mathcal{X}^{2})-\frac{1}{n}\left(\mathrm{tr}\mathcal{X}\right)^{2}$
@@
$\left\lvert\mathscr{X}\right\rvert^{2}=\left(\mathcal{X}_{i}^{j}-\frac{\mathrm{tr}\mathcal{X}}{n}\delta_{i}^{j}\right)\left(\mathcal{X}_{i}^{j}-\frac{\mathrm{tr}\mathcal{X}}{n}\delta_{i}^{j}\right)=\mathrm{tr}\left(\mathcal{X}\mathcal{X}^{T}\right)-\frac{1}{n}\left(\mathrm{tr}\mathcal{X}\right)^{2}.$
@@
Remark 3.4.
From the proof of the above Lemma 3.3, it is easy to see that for the tensor $\mathscr{X}$ difined as before, where $u$ is a $C^{3}$-smooth positive function with $|\nabla u|\neq 0$ on $\Omega$ (in other words, we do not need assume that $u$ is a solution to the equation (2.1)), there holds pointwisely true the same inequality
@@
We compute from (3.2) that
@@
It is worthy to point out that the above lemma holds true for any $u\in C^{3}(\Omega)$ with $\nabla u\neq 0$ a.e. on $\Omega$ and one does not need to assume that $u$ satisfies the equation (2.1). Applying this lemma we can derive the following fundamental and important pointwise identity.
@@
$\begin{split}&\operatorname{div}\left(u^{a}\left(\mathscr{X}(\mathbf{X})\right)-\dfrac{n(p-1)a+(n-1)pb}{np}u^{a-1}\left\langle\mathbf{X},\nabla u\right\rangle\mathbf{X}\right)\\ =&u^{a}\left(\mathrm{tr}\left(\mathscr{X}^{2}\right)+\mathrm{Ric}\left(\mathbf{X},\mathbf{X}\right)\right)\\ &+\left(\dfrac{(p-1)a}{p}\left(1-\dfrac{(n-p)a}{(n-1)p}\right)-\dfrac{n-1}{n}\left(b+\dfrac{n(p-1)a}{(n-1)p}\right)^{2}\right)u^{a+2b-2}\left\lvert\nabla u\right\rvert^{2p}\\ &+\dfrac{n-1}{n}\left(\dfrac{((n+1)p-n)a}{(n-1)p}f(u)-uf^{\prime}(u)\right)u^{a+2b-1}\left\lvert\nabla u\right\rvert^{p}.\end{split}$ (3.8)
@@
$\begin{split}&\operatorname{div}\left(u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\left(\mathscr{X}(\mathbf{X})\right)\right)\\ =\,&u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\left(\mathrm{tr}\left(\mathscr{X}^{2}\right)+\mathrm{Ric}\left(\mathbf{X},\mathbf{X}\right)\right)\\ &+\dfrac{(n-1)(p-1)\alpha}{(n+1)p-n}\left(1-\dfrac{(n-p)\alpha}{(n+1)p-n}\right)u^{-\frac{((n+1)p-2n)\alpha}{(n+1)p-n}-2}\left\lvert\nabla u\right\rvert^{2p}\\ &+\dfrac{n-1}{n}\left(\alpha f(u)-uf^{\prime}(u)\right)u^{-\frac{((n+1)p-2n)\alpha}{(n+1)p-n}-1}\left\lvert\nabla u\right\rvert^{p}.\end{split}$ (3.9)
@@
$=$ $u^{a}\left(\mathrm{tr}\left(\mathscr{X}^{2}\right)+\mathrm{Ric}\left(\mathbf{X},\mathbf{X}\right)\right)$
@@
$=\,$ $\operatorname{div}\left(u^{a+2b}\Delta_{p}u\mathbf{U}\right)-\left\langle\nabla\left(u^{a+2b}\Delta_{p}u\right),\mathbf{U}\right\rangle$
@@
$=\,$ $u^{a}\left(\mathrm{tr}\left(\mathscr{X}^{2}\right)+\mathrm{Ric}\left(\mathbf{X},\mathbf{X}\right)\right)$
@@
we get the desired formula (3.9) and complete the proof formally. By 3.2, we know that $\mathrm{tr}\left(\mathcal{X}^{2}\right)\in L^{1}_{loc}$. It follows that every term on the right hand side of 3.6 lies in $L^{1}_{loc}$. ∎
In the end of this section, we collect some useful inequalities in geometric analysis. Let $M^{n}$ be a Riemannian manifold with Ricci curvature bounded from below by some nonnegative constant $(n-1)\kappa$. The first one is the following Saloff-Coste’s Dirichlet Sobolev inequality: for every $\psi\in C_{0}^{\infty}\left(B_{R}\right)$
[…]